two lady bugs sit on a rotating disk without slipping. lady bug 1 is half away between the rotation axis and lady bug 2. the linear speed v, of ladybug 1 is

a. half that of ladybug 2

b. the same as ladybug 2

c. twice that of ladybug 2

d. impossible to determine from the information given
Answer #1

the correct option is a) half that of ladybug 2

we know the formula for linear speed is

v=r, so we have

v1=r1 and v2=r2

here r1 is half of total radius so we have r1=r/2 and r2=r

so we have the ratio

v1/v2=(r/2/r)=1/2

v1=1/2 *v2

so linear speed of 1 is half of ladybug 2